IF 4. FIND THE MASS OF A CONICAL FUNNEL Z= 7x+y Oz<4 THE DENSITY PER UNIT AREA IS p=8-2.
Answer #1

I have assumed the cone to be open at top. AND since density is given in per area terms, take the lateral surface area of the cone calculating the mass.( If u want to change my assumption, let me know.)

+ z = x² + y² z O<Z24 z = x²+4² P = 8-2 en mass Area Zo Area at since density is in Mass/ Area Take the lateral surface heigh